A cross-country flight traveling with the wind covers 3000 miles in 5 hours. The same plane changes direction and travels against the wind, covering 1500 miles in 3 hrs. What are the wind speed and the speed of the plane in still air?
Answer by jorel1380(3719)

3000/5=600 mph
1500/3=500 mph
600+500/2=550 mph
550+x=600
550-x=500
2x=100
x=50
The speed of the plane in still air is 550 mph. The wind speed is 50 mph..